OpenStudy (anonymous):

A company will need $35,000 in 8 years for a new addition. To meet this goal, the company deposits money in an account today that pays 6% annual interest compounded quarterly. Find the amount that should be invested to total $35,000 in 8 years. I know the formula is A=P(1+.06/4)^8 A=35000(1+.015)^8 A=35000(1.015)^8 A=35000(1.1264925866) A=39427 (rounded to nearest dollar) Is this correct?

OpenStudy (mertsj):

The formula is: \[A=P(1+\frac{r}{n})^{nt}\]

OpenStudy (anonymous):

That's what I did I thought A=35000(1+6%/4)^8(years)

OpenStudy (anonymous):

the 6%/4(quarterly)

OpenStudy (mertsj):

What is n?

OpenStudy (anonymous):

yeah

OpenStudy (mertsj):

yeah is not a valid replacement for n

OpenStudy (anonymous):

the rate being 6% annually and the n being quarterly for 4 quarters in a year

OpenStudy (mertsj):

Are you saying that n is 4?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Then what is nt?

OpenStudy (anonymous):

then the exponent is 8 for 8 years right?

OpenStudy (mertsj):

It's a fairly simple concept. n = 4, t = 8, nt=32. The exponent is nt. Therefore the exponent is 32

OpenStudy (anonymous):

oh...Ok let me figure this out LOL

OpenStudy (anonymous):

$56,361 ?

OpenStudy (mertsj):

Did you read the problem?

OpenStudy (anonymous):

yeah it said to round to the nearest dollar

OpenStudy (mertsj):

Do you really think it makes sense to say they must invest more now then they will need in 8 years? Are they going to lose money on this investment or what?

OpenStudy (anonymous):

Ok now I am really confused...A=35000(1+0.06/4)^32 This is correct right?

OpenStudy (mertsj):

The AMOUNT they need in 8 years is 35000. What PRINCIPAL should they invest now so that the AMOUNT they will have in 8 years is 35000?

OpenStudy (anonymous):

GRR I so dumb 35000=P(1+.06/4)^32

OpenStudy (anonymous):

I wasn't even picking up on that..

OpenStudy (anonymous):

Ok next would be 35000=P(1.015)^32 then 35000=P(1.6103243202) This is where I am getting stuck

OpenStudy (mertsj):

Solve this equation: 24=6x

OpenStudy (anonymous):

6*4=24

OpenStudy (mertsj):

How did you get that?

OpenStudy (anonymous):

multiplication

OpenStudy (anonymous):

$21,734

OpenStudy (mertsj):

The correct answer (which you would know if you had followed the instructions of your algebra I instructor to show your work) is "divide both sides by 6.

OpenStudy (anonymous):

divide 35000 by 1.6103243202

OpenStudy (anonymous):

Im not taking Algebra I am in MATH 156 how to teach elementary math this was the only problem I was stuck on =o)

OpenStudy (anonymous):

I realized that I after you pointed out that I put the 35000 in the wrong spot...Thanks for the help!!!

OpenStudy (mertsj):

Good for you and you are welcome.